OpenStudy (anonymous):

12.5The vertices of a tetrahedron are located at the points (1,2,-1), (3,1,- 2) , (-1,-2,0) and (2,1,5) . Find the altitude of tetrahedron measured from the vertex at (2,1,5)

OpenStudy (amistre64):

so the distance of a point from a plane

OpenStudy (amistre64):

point of (2,1,5) plane of: (1,2,-1), (3,1, -2) , (-1,-2,0) -1-2 1 -1-2 1 -1-2 1 ------------------------- <0,0,0> <2,-1,-1> <-2,-4,1> cross the 2 nonzero vectors for a normal x 2 -2 x = -1-2 = -3 y -1 -2 -y = 2-2 = 0 z -1 1 z = -4-2 = -6 our normal to the plane is: <3,0,2>, and anchor it to our point as a line eq: x= 2+3t y= 1 z= 5+2t when this line intersects the plane; we can define the point that is at the perp under the point

OpenStudy (amistre64):

define our plane by the normal vector and any point in it; and make sure our normal isnt messed up by forgetting how to divide by -3 lol <-3,0,-6> ----------- = <1,0,2> so change x to x = 2+t -3

OpenStudy (amistre64):

1(x-1)+0(y-2)+2(z+1)=0 2+t -1 +10+4t +2 = 0 t = 3/5

OpenStudy (amistre64):

insert that into the line equation to determine the points to distance x= 10+ 3/5 = 13/5 y= 1 = 5/5 z= 25+6/5 = 31/5 those look lovely :)

OpenStudy (amistre64):

2,1,5 = 10/5, 5/5, 25/5 (13/5, 5/5, 31/5) -(10/5, 5/5, 25/5) ----------------- 3/5 , 0 , 6/5 finish out the distance process and we get sqrt(9+36) ---------- 5

OpenStudy (amistre64):

of course dbl chk the work to make sure I didnt do any basic mathing errors :)

OpenStudy (anonymous):

Thanks man, i,m right on it.

OpenStudy (amistre64):

my t value is off ....

OpenStudy (amistre64):

i totally ignored the +10 :) 1(2+t-1)+0()+2(5+2t+1)=0 (1+12)2 +t(1+4) = 0 t = -13/5

OpenStudy (amistre64):

x= 10/5 - 13/5 = -3/5 y= 5/5 = 5/5 z= 25/5 -26/5 = -1/5

OpenStudy (amistre64):

(-3/5, 5/5, -1/5) -(10/5, 5/5, 25/5) ----------------- -13/5 , 0 , 26/5 sqrt(159+676) ------------- ; maybe :) 5

OpenStudy (amistre64):

my fingers hate me :) 13^2 = 169 not 159

OpenStudy (amistre64):

13/sqrt(5) i recall it being something similar to the t value